# Multi-view Diffusion for 3D Generation

### v3

@@ 3.2 Text-to-3D generation
We also use the point lighting (Poole et al., 2023) and soft shading (Lin et al., 2023a) to regularize the geometry. For regularization loss, we only use the orientation loss proposed by (Poole et al., 2023). We note that both of these two techniques only help to smooth the geometry and have little effect on the content in our case. We do not use any sparsity loss to force the separation between foreground the background but instead achieve it by replacing background with random colors. It is also worth noting that MVDream is theoretically combinable with other SDS variants such as SJC (Wang et al., 2023a) and VSD(Wang et al., 2023c), but this is out of the scope of this paper.
@@ 4.1 Multi-view Image Generation
Validation set
N/A $12.90\pm 0.66$ $30.12\pm 3.15$
Multi-view Diffusion (3D data only)
$40.38$ $12.33\pm 0.63$ $29.69\pm 3.36$
Multi-view Diffusion (3D + LAION 2D data)
$39.04$ $12.97\pm 0.60$ $30.38\pm 3.50$
@@
${}^{\circ}$
${}^{\circ}$
${}^{\circ}$
${}^{\circ}$
${}^{\circ}$
${}^{\circ}$
${}^{\circ}$
${}^{\circ}$
@@
5 Discussion and Conclusion
@@ Appendix A Appendix
A.1 Comparison with Image-to-3D Method
In Fig. (10), we compare our results with an state-of-the-art image-to-3D method, namely Zero123-XL (Liu et al., 2023). In particular, we first generate an image using the same prompt with SDXL (Podell et al., 2023), segment the object to remove background, and then provide the image as input to the image-to-3D system using Zero123-XL in threestudio. We notice two drawbacks of such a pipeline compared to our text-to-multi-view diffusion model:
  1. 1.
    Zero123-XL is trained on 3D datasets only, whose objects are mostly aligned with simple poses (e.g. T-pose for characters). So the image-to-3D results tend to be distorted if the input image has complicated poses and view angles that are different from 3D dataset. Specifically, when the object is geometrically spanned in the direction of the depth, the generated 3D tends to be too flat. This can be seen from the examples of eagle.
  2. 2.
    Zero123-XL can hardly generate geometrically consistent and richly textured novel-view images, and therefore the score distilled objects tend to have less details, yielding a smooth, blurred and relatively uniform appearance on the other sides different from the input view, as shown with the example of bulldog.
Zero123-XL Refer to caption
Ours Refer to caption
(a) a bald eagle carved out of wood
Zero123-XL Refer to caption
Ours Refer to caption
(b) a bulldog wearing a black pirate hat
Zero123-XL Refer to caption
Ours Refer to caption
(c) beautiful, intricate butterfly
1 view
2 view
4 view (proposed)
an astronaut riding a horse
baby yoda in the style of Mormookiee
3D data only
3D + 2D (LAION) data
@@ A.1 Comparison with Image-to-3D Method
A.2 Effect of Number of Views
A.3 Training with 2D data
@@
A.4 Training with random camera
In the main paper, our model is trained for 4 orthogonal views with the same elevation angle. Thus, a question arises: can the method generalize to scenarios involving random camera views rather than orthogonal ones? We conduct such an experiment and found that the proposed 3D self-attention network can indeed generate random multi-view images if trained on such data. In spite of worse quality compared to the orthogonal model, an interesting finding is that such a random-view model able to generate a much larger number of views during inference, as shown in Fig. (13). In our experiments, we found it able to generate 64 views and potentially more, even though it is trained on 4 views.
Figure 13: Multi-view diffusion model trained on random views is able to generate more a larger number of views than training (40 views for example here).
A.5 $\mathbf{x}_{0}$ Reconstruction Loss for SDS
@@ A.6 Implementation Details
A.6.1 Data Preparation and Diffusion Model Training

### v4

@@ 3.2 Text-to-3D generation
In contrast to prior studies that incorporates many regularization techniques Poole et al. (2023); Lin et al. (2023a); Wang et al. (2023c), we found that such a simple reconstruction loss in Eq. (2) loss is sufficient for 3D generation from scratch, indicating our multi-view diffusion as a strong 3D prior. This being said, we still consider to use the point lighting (Poole et al., 2023) and soft shading (Lin et al., 2023a) to improve the geometry of the output. For regularization loss, we only use the orientation loss proposed by (Poole et al., 2023). We note that both of these two techniques only help to smooth the geometry and makes little different to the generated content in our case. We do not use any sparsity loss to force the separation between foreground the background but instead achieve it by replacing background with random colors. It is also worth noting that MVDream is theoretically combinable with other SDS variants such as SJC (Wang et al., 2023a) and VSD(Wang et al., 2023c), but this is out of the scope of this paper.
@@ 4.1 Multi-view Image Generation
Validation set N/A $12.90\pm 0.66$ $30.12\pm 3.15$
Multi-view Diffusion (3D data only) $40.38$ $12.33\pm 0.63$ $29.69\pm 3.36$
Multi-view Diffusion (3D + LAION 2D data) $39.04$ $12.97\pm 0.60$ $30.38\pm 3.50$
@@
0 45 90 135 0 45 90 135
@@
Figure 7: User study.
1 view 2 view 4 view (proposed)
Refer to caption Refer to caption Refer to caption
an astronaut riding a horse
Refer to caption Refer to caption Refer to caption
baby yoda in the style of Mormookiee
@@
5 Discussion
Generalizability to More Views
In the main paper, our model is trained for 4 orthogonal views with the same elevation angle. Thus, a question arises: can the method generalize to scenarios involving more camera views that are not orthogonal to each other? To answer the question, we conduct a simple extension experiment and found that the proposed 3D self-attention network can indeed generate random multi-view images if trained on such data. In spite of worse quality compared to the orthogonal model, an interesting finding is that such a random-view model able to generate a much larger number of views during inference, as shown in Fig. (11). In our experiments, we found it able to generate 64 views and potentially more, even though it is trained on 4 views. Although this many views are unnecessary for the proposed multi-view SDS, we believe this finding could be beneficial for future research on multi-view image generation.
Figure 11: Multi-view diffusion model trained on random views is able to generate a larger number of views than training (In this example we used 4 views for training and 40 views for inference).
Comparison with Image-to-3D Method
In Fig. (12), we compare our results with an state-of-the-art image-to-3D method, namely Zero123-XL (Liu et al., 2023). In particular, we first generate an image using the same prompt with SDXL (Podell et al., 2023), segment the object to remove background, and then provide the image as input to the image-to-3D system using Zero123-XL in threestudio. We notice two drawbacks of such a pipeline compared to our text-to-multi-view diffusion model: (1) Zero123-XL is trained on 3D datasets only, whose objects are mostly aligned with simple poses (e.g. T-pose for characters). So the results tend to be distorted if the input image has complicated poses. (2) Zero123-XL can hardly generate geometrically consistent novel-view images, and therefore the score distilled objects tend to have less details, yielding a blurred appearance on the other sides from the input view.
Zero123-XL Refer to caption
Ours Refer to caption
(a) a bald eagle carved out of wood
Zero123-XL Refer to caption
Ours Refer to caption
(b) a bulldog wearing a black pirate hat
@@ Appendix A Appendix
3D data only 3D + 2D (LAION) data
Refer to caption Refer to caption
@@
A.1 2D Data
@@
A.2 $\mathbf{x}_{0}$ Reconstruction Loss for SDS
@@ A.3 Implementation Details
A.3.1 Data Preparation and Diffusion Model Training